OpenStudy (anonymous):

Thrifty Supply Store normally sells candy bars for $0.59. Today they are running a special 2 candy bars for $0.99. How much would you save if you bought 10 candy bars at the sale price rather than at the regular price?

OpenStudy (anonymous):

normal rate=0.59*10 offered rate= (0.99/2)*10 now gain=normal rate- offered rate. . .